# Broadcast in Almost Mixing Time

### v2

@@ Abstract
Our primary contribution is a randomized algorithm for networks with expander topology, which are widely used in practice for building scalable and robust distributed systems. The algorithm succeeds with high probability and achieves a round complexity that is optimal up to a factor of the network’s mixing time and polylogarithmic terms. It leverages a multi-COBRA primitive, which uses multiple branching random walks running in parallel. To the best of our knowledge, this approach has not been applied in distributed algorithms before. A crucial aspect of our method is the use of these branching random walks to construct an optimal (up to a polylogarithmic factor) tree packing of a random graph, which is then used for efficient broadcasting. This result is of independent interest.
We also prove the problem to be NP-hard in a centralized setting and provide insights into why straightforward lower bounds for general graphs, namely graph diameter and $\frac{|M|}{\textit{minCut}}$, cannot be tight.
@@
Distributed algorithms, Expander Graphs, Random graphs, Broadcasting, Branching random walks, Tree packing, CONGEST model.
@@ 1 Introduction
Expander graphs possess a unique combination of highly desirable properties for networked systems, including low diameter, high connectivity, rapid mixing, and strong resilience to node failures. These characteristics make them ideal topologies for ensuring fast information dissemination, robust communication, and efficient resource discovery in distributed environments. As a result, overlay network topologies of major distributed systems are often designed to emulate the properties of expander graphs [28, 43, 34, 50, 14, 4, 51, 46]. This practical adoption of high-expansion structures further reinforces the motivation to study and design algorithms specifically optimized for expander graphs, aiming to leverage their theoretical advantages in real-world distributed systems.
One of the fundamental communication tasks in such distributed systems is to efficiently broadcast multiple messages. In blockchains, this could be the dissemination of transactions, while in peer-to-peer networks such as Torrent, it may involve distributing chunks of a large file. The design of robust and efficient broadcast algorithms is critical for maintaining the performance and scalability of these systems.
In this work, we present an algorithm that addresses the problem of broadcasting multiple messages in distributed systems with an expander graph topology. Our approach broadcasts information nearly optimally, with an overhead proportional to the network’s mixing time and a small polylogarithmic factor while imposing no requirements on the underlying network topology beyond its expansion properties. This generality ensures that our approach is applicable to a wide range of real-world systems and highlights its potential as a versatile tool for scalable and efficient information dissemination. Let us now formally define the setting.
@@ 1.1 Model and Problem
The CONGEST model [47] is defined as follows. The network is modeled as a graph with $n$ nodes, where execution progresses in synchronous rounds. In each round, a node can send a message of size $O(\log n)$ bits to each of its neighbors. Nodes do not have prior knowledge of the network topology but are assumed to have unique identifiers that fit within $O(\log n)$ bits.
@@
A dedicated source node $s$ possesses a set $M$ of messages, where each message $m\in M$ has a size of $O(\log n)$ bits. The objective is to ensure that every node in the network learns all messages in $M$.
As pointed out by Ghaffari [23], the problem suggests an $\Omega(D+k)$ round complexity lower bound, where $D$ is the diameter of the graph and $k$ is the number of messages $|M|$. For example, consider a path graph with $s$ as its first node. Any algorithm would require at least $D+k-1$ rounds to transmit all messages to the last node. However, this bound is only existential, meaning there exists a graph for which $\Omega(D+k)$ rounds are needed. In contrast, consider a complete graph with $k=n$. Here, broadcasting can be completed in $2$ rounds, which is significantly better than the $\Omega(k)=\Omega(n)$ bound suggested by the path graph example. Our paper presents an algorithm that achieves universal optimality [22] on expander graphs. Specifically, for any expander graph, our algorithm completes the multi-message broadcast in a number of rounds that is within a small overhead of the best possible for that specific graph. Before we state our results formally, we need to introduce helpful terminology.
@@ 1.2 Preliminaries and Notation
One of the key graph-theoretical components in our approach is tree packing. A tree packing of a graph $G$ is a collection of spanning subtrees of $G$. The tree packing is characterized by three parameters: (1) its size $S$, i.e., the number of trees, (2) its height $H$, i.e., the maximal height of a tree, and (3) its weight $W$, i.e., the maximal number of trees sharing a single edge.
The present work focuses specifically on two graph families, namely Erdős–Rényi graphs and expanders. An Erdős–Rényi graph $G(n,p)$ is a graph on $n$ vertices where each edge exists independently from others with probability $p$ [18]. We refer to a graph as an expander if it has small (polylogarithmic) mixing time [^1] An alternative way is to say that expander graphs are those that feature an inverse polylogarithmic conductance; these two notions are equivalent. See Appendix A for a definition of conductance..
For our purposes, it will be convenient to define the mixing time of an undirected graph by reconsidering it as being bidirected, that is, with each undirected edge $(u,v)$ replaced with $(u,v)$ and $(v,u)$. The mixing time $\tau_{mix}$ of a bidirected graph is defined as follows. Consider a lazy random process that starts at an arbitrary edge of the graph. At each step, with probability $\frac{1}{2}$, the process remains at the current edge, and with probability $\frac{1}{2}$, it transitions to a uniformly random adjacent edge (a directed edge $e_{2}$ is adjacent to a directed edge $e_{1}$ if they are of the form $e_{1}=(u,v)$ and $e_{2}=(v,w)$). It is known that this process admits a stationary distribution $\pi$, which is uniform over all edges. Moreover, regardless of the starting edge, the distribution $D_{t}$ of the walk after $t$ steps converges to $\pi$. We define $\tau_{mix}$ as the smallest $t$ for which $D_{t}$ is inverse-polynomially close to $\pi$. For a formal definition, please refer to Appendix A. We are now ready to formally state our results.
@@ 1.3 Our Contribution
Our main result is an algorithm to solve the multi-message broadcast problem with only an overhead of $\tilde{O}(\tau_{mix})$.
@@
There exists a randomized distributed algorithm that for any graph $G$ solves the multi-message broadcast problem in $O(\log^{3}n\cdot\tau_{mix}\cdot\mathrm{OPT})$ rounds with high probability, where $\tau_{mix}$ is the mixing time of the network graph and $\mathrm{OPT}$ is the optimal round complexity for the given problem instance.
We design the algorithm for a general expander by building on the algorithm for random graphs, which achieves near-optimal performance in an important special case when the network is modeled as an Erdős–Rényi graph $G(n,p)$.
Throughout the paper, let $C_{p}$ denote a sufficiently large constant. [^2] It suffices to take $C_{p}=2700$. It was not a concern for the present work to optimize this constant.
@@
Note that the round complexity of $O(\log^{2}n+\log n\cdot\frac{k}{\delta(G)})$ is optimal up to an additive $\log^{2}n$ term and a multiplicative $\log n$ factor since $\frac{k}{\delta(G)}$ is a straightforward lower bound. Also, the condition $p=\Omega(\frac{\log n}{n})$ is necessary, since for $p\leq\frac{\log n}{n}$, there is a constant probability that the graph is disconnected [18].
@@
For an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, there exists a distributed randomized algorithm that produces a tree packing of size $\delta(G)$, height $O(\log n)$ and weight $O(\log n)$ w.h.p.
@@
Finally, to map the terrain of the problem, we prove that it is NP-hard in the centralized case to achieve the optimal number of rounds for general graphs. Also, we provide an instance of a problem where $D(G)$ and $\frac{k}{\textit{minCut}(G)}$ – two straightforward lower bounds for round complexity – are both $O(1)$, while the optimal round complexity is $\Omega(\sqrt{k})$.
@@ 2 Related Work
Previous Work. The first work to address universal optimality for the multi-message broadcast problem in the CONGEST model was by Ghaffari [23]. In that paper, the algorithm consists of two phases: (1) constructing a tree packing, and (2) performing the broadcast using the constructed tree packing. With a tree packing of height $H$, size $S$ and weight $W$, one can complete a multi-message broadcast in $O((H+\frac{k}{S})\cdot W)$ rounds by splitting messages uniformly across the trees and propagating them sequentially within each tree. However, the limitation of [23] is that constructing the tree packing requires $\tilde{\Omega}(D+k)$ rounds, preventing the approach from achieving universal optimality.
@@
Routing. One fundamental information-dissemination problem in distributed computing is routing, where the goal is to deliver a set of messages from source nodes to their respective destination nodes. Unlike broadcast, which involves sending the same message(s) to all nodes, routing requires sending individual message(s) for each source-destination pair.
@@
[…]
@@
Broadcast in the Presence of Churn. To account for the instability of connections in distributed systems, researchers often consider models where the graph topology undergoes modifications while the algorithm is running. Kuhn et al. [37] introduced a model where, in each round, the communication graph is chosen adversarially, and achieved a round complexity of $O(nk)$ for broadcasting $k$ messages in this setting. As a follow-up, Dutta et al. [16] established a lower bound of $\tilde{\Omega}(nk)$ for this problem, providing a precise sequence of adversarially constructed graphs that makes the problem difficult. However, Dinitz et al. [13] showed that this lower bound is “fragile.” Specifically, they demonstrated that the hard case is carefully constructed to prevent algorithms from succeeding. They did so by introducing the concept of $l$-smoothing to the adversarial sequence of graphs, where each graph in the sequence additionally gets $l$ edges added at random. Surprisingly, even with $l=1$, an algorithm that broadcasts a random message at possession to its neighbors in each round achieves a round complexity of $\tilde{O}(n+k^{3})$ w.h.p., effectively breaking the lower bound of [16].
Tree Packing. The problem of tree packing has been extensively studied, as summarized in the survey by Palmer [45]. Foundational results in this area include those by Tutte [53] and Nash-Williams [44], who demonstrated that an undirected graph with edge connectivity $\lambda$ contains a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$. Edmond [17] extended this result to directed graphs, showing that such graphs always contain $\lambda$ pairwise edge-disjoint spanning trees rooted at a sender $s\in V$, where $\lambda$ is the minimum number of edges that must be removed to make some node unreachable from $s$. However, these results do not address the height of the tree packing.
Chuzhoy et al. [11] tackled the challenge of finding tree packings with small height. They presented a randomized algorithm that, given an undirected graph with edge connectivity $\lambda$ and diameter $D$, outputs with high probability a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$, weight $2$, and height $O((101k\log n)^{D})$.
Tree packing on random graphs was studied by Gao et al. [21], who showed that asymptotically almost surely, the size of a spanning tree packing of weight one for a Erdős–Rényi graph $G(n,p)$ is $\min\left\{\delta(G),\frac{|E(G)|}{n-1}\right\}$, which corresponds to two straightforward upper bounds.
@@
Branching Random Walks in Networks. The cover time of a random walk [39] on a graph is the time needed for a walk to visit each node at least once. Unfortunately, the expected value of this quantity is $\Omega(n\log n)$ even for a clique, making this primitive less useful in practical applications. Consequently, several attempts have been made to accelerate the cover time. Alon et al. [2] proposed initiating multiple random walks from a single source. Subsequent work by Elsässer and Sauerwald refined their bounds, demonstrating that $r$ random walks can yield a speed-up of $r$ times for many graph classes. Variations of multiple random walks have been applied in the CONGEST model to approximate the mixing time [42], perform leader election [38, 27], and evaluate network conductance [19, 5].
@@ 3 Algorithm Overview
In this section, we provide a high-level overview of our algorithm, which consists of two major parts. First, we embed a virtual Erdős–Rényi graph $G(n,p)$ atop the underlying (physical or overlay) network $H$, and then we solve the problem on $G$. We explain the embedding procedure in Section 3.1, and from that point on, we focus solely on solving multi-message broadcast on an Erdős–Rényi graph. In Section 3.2, we overview COBRA - the main building block of our algorithm for random graphs, and in Section 3.3 we describe the algorithm itself.
@@ 3.1 Embedding a Random Graph
Our construction will guarantee that each round of communication in $G$ can be simulated in $O(\tau_{mix}(H)\cdot\log n)$ rounds in $H$ and that $\delta(G)$ will be close to $\delta(H)$. This gives us the following “lifting”.
@@
Consider a graph $H$ with $|V(H)|=n$. Assume there exists an algorithm that solves multi-message broadcast on an Erdős–Rényi graph $G(n^{\prime},\frac{\delta(H)}{n^{\prime}})$ in $R$ rounds w.h.p. for some $n^{\prime}=poly(n)$. Then there exists an algorithm that solves multi-message broadcast on $H$ in $O(R\cdot\tau_{mix}\cdot\log n+\tau_{mix}\cdot\log^{2}n)$ rounds w.h.p.
This round complexity consists of three terms: $O(R\cdot\tau_{mix}\cdot\log n)$ for simulating an algorithm, $O(\tau_{mix}\cdot\log^{2}n)$ rounds for constructing an embedding knowing $\tau_{mix}$, and $O(\tau_{mix}\cdot\log^{2}n)$ for estimating the $\tau_{mix}$. Below, we describe both the embedding and estimation parts.
Embedding Construction. We now outline how to construct the embedding of an Erdős–Rényi graph atop the expander $H$, assuming nodes have an estimate of $\tau_{mix}$.
@@
Once paths are established, communication along $(u,v)\in E(G)$ is simulated by routing a message along the corresponding walk path in $H$, which, by Lemma 3.2, can be done in $O(\tau_{mix}\cdot\log n)$ rounds.
We remark that the process described above is equivalent to sampling with replacement $\delta$ neighbors for each node $v\in V(G)$ independently (almost) uniformly at random. While this is not a canonical definition of an Erdős–Rényi graph, the resulting distribution is equivalent, modulo a negligible inverse-polynomially small probability.
On Distributed Estimation of Mixing Time. Note that in order to implement this simulation, nodes do not need prior knowledge of $\tau_{mix}$. Instead, they can estimate the mixing time of the network using the decentralized algorithm of Kempe and McSherry [36], which runs in $O(\log^{2}n\cdot\tau_{mix})$ rounds (without prior knowledge of $\tau_{mix}$). Their algorithm can be used to estimate the second principal eigenvalue $\lambda$ of the transition matrix, which relates to the mixing time via the following inequality:
@@
Given a graph on $n$ nodes with mixing time $\tau_{mix}$ and second principle eigenvalue $\lambda$, it holds that
@@
Knowing $\tau_{mix}(H)$ and $\delta(H)$, compute an embedding as described above. Make the source node for a multi-message broadcast in $G$ any node in $G$ that is simulated by the source of an original problem instance in $H$. Solve the problem in $G$ in $R$ rounds, simulating each round of $G$ in $O(\tau_{mix}(H)\cdot\log n)$ rounds in $H$.

### v3

@@ Abstract
Our primary contribution is a randomized algorithm for networks with expander topology. The algorithm succeeds with high probability and achieves a round complexity that is optimal up to a factor of the network’s mixing time and polylogarithmic terms. It leverages a multi-COBRA primitive, which uses multiple branching random walks running in parallel. A crucial aspect of our method is the use of these branching random walks to construct an optimal (up to a polylogarithmic factor) tree packing of a random graph, which is then used for efficient broadcasting.
We also prove the problem to be NP-hard in a centralized setting and provide insights into why lower bounds that can be matched in expanders, namely graph diameter and $\frac{|M|}{\textit{minCut}}$, cannot be tight in general graphs.
@@
Distributed algorithms, Expander Graphs, Random graphs, Broadcast, Branching random walks, Tree packing, CONGEST model.
@@ 1 Introduction
The CONGEST model [42] was originally introduced to study computer networks with limited communication channels. Though later, despite its simplicity, it has given rise to a wide range of theoretical challenges. Progress on those has not only allowed for faster distributed algorithms, but also led to a deeper understanding of fundamental graph abstractions (see, e.g., [23, 22, 15, 44, 6]).
The problem we address in this work fits precisely into this tradition. We study the broadcast of multiple messages, where a single source node holds a collection of messages that must be disseminated so that every node in the network receives them as quickly as possible. In real-world systems, this setting naturally models the distribution of data chunks in peer-to-peer file sharing or block chunks dissemination in blockchain protocols. Theoretically, however, the problem reduces to the tree packing in the underlying network graph - a concept still not fully understood.
We show that this problem admits a fast solution in expander networks. In particular, we present an algorithm that broadcasts multiple messages with near-optimal round complexity: the overhead depends only on the network’s mixing time, up to a small polylogarithmic factor. The key technical contribution enabling this result is a nearly-optimal distributed tree packing procedure for a random graph, which serves as the structural backbone of our dissemination strategy and may be of independent interest beyond this application.
We now proceed to define the setting formally.
@@ 1.1 Model and Problem
The CONGEST model is defined as follows. The network is modeled as a graph with $n$ nodes, where execution progresses in synchronous rounds. In each round, a node can send a message of size $O(\log n)$ bits to each of its neighbors. Importantly, nodes do not have prior knowledge of the network topology.
@@
A dedicated source node $s$ possesses a set $M$ of $k$ messages, where each message $m\in M$ has a size of $O(\log n)$ bits. The objective is to ensure that every node in the network learns all messages in $M$.
As pointed out by Ghaffari [21], the problem suggests an $\Omega(D+k)$ round complexity lower bound, where $D$ is the diameter of the graph. For example, consider a path graph with $s$ as its first node. Any algorithm would require at least D+k1D+k-1 rounds to transmit all messages to the last node. However, this bound is only existential, meaning there exists a graph for which $\Omega(D+k)$ rounds are needed. In contrast, consider a complete graph with $k=n$. Here, broadcasting can be completed in two rounds: in the first round, $s$ sends the $i$-th message to the node $i$, and in the second round, each node broadcasts the message it received in round $1$. This is significantly better than the $\Omega(k)=\Omega(n)$ bound suggested by the path graph example. These contrasting cases highlight the importance of algorithms that adapt to the underlying topology. Our paper presents such an algorithm, achieving universal optimality [20] on expander graphs. Specifically, it completes the multi-message broadcast on every expander $G$ in a number of rounds within a small overhead of the best possible for $G$. Before stating our results formally, we introduce some necessary terminology.
@@ 1.2 Preliminaries and Notation
One of the key graph-theoretical components in our approach is tree packing. A tree packing of a graph $G$ is a collection of spanning subtrees of $G$. The tree packing is characterized by three parameters: (1) its size $S$, i.e., the number of trees, (2) its diameter $H$, i.e., the maximal diameter of a tree, and (3) its weight $W$, i.e., the maximal number of trees sharing a single edge.
The present work focuses specifically on two graph families, namely Erdős–Rényi graphs and expanders. An Erdős–Rényi graph $G(n,p)$ is a graph on $n$ vertices where each edge exists independently from others with probability $p$ [16]. Throughout the paper, let $C_{p}$ denote a sufficiently large constant [^1] It suffices to take $C_{p}=2700$. It was not a concern for the present work to optimize this constant. such that $G(n,p)$ is connected w.h.p. for $p\geq\frac{C_{p}\log n}{n}$. The condition $p=\Omega(\frac{\log n}{n})$ is necessary, since for $p\leq\frac{\log n}{n}$, there is a constant probability that the graph is disconnected [16].
We refer to a graph as an expander if it has small (polylogarithmic) mixing time [^2] An alternative way is to say that expander graphs are those that feature an inverse polylogarithmic conductance; these two notions are equivalent. See Appendix A for a definition of conductance.. For our purposes, it will be convenient to define the mixing time of an undirected graph by reconsidering it as being bidirected, that is, with each undirected edge $(u,v)$ replaced with $(u,v)$ and $(v,u)$. The mixing time $\tau_{mix}$ of a bidirected graph is defined as follows. Consider a lazy random process that starts at an arbitrary edge of the graph. At each step, with probability $\frac{1}{2}$, the process remains at the current edge, and with probability $\frac{1}{2}$, it transitions to a uniformly random adjacent edge (a directed edge $e_{2}$ is adjacent to a directed edge $e_{1}$ if they are of the form $e_{1}=(u,v)$ and $e_{2}=(v,w)$). It is known that this process admits a stationary distribution $\pi$, which is uniform over all edges. Moreover, regardless of the starting edge, the distribution $D_{t}$ of the walk after $t$ steps converges to $\pi$. We define $\tau_{mix}$ as the smallest $t$ for which $D_{t}$ is inversely polynomial close to $\pi$. For a formal definition, please refer to Appendix A. We are now ready to formally state our results.
@@ 1.3 Our Contribution
Our main result is an algorithm to solve the multi-message broadcast problem with only an overhead of $\tilde{O}(\tau_{mix})$. We highlight that our algorithm can be run on any graph, but it will only be efficient on graphs with small $\tau_{mix}$, i.e., expanders.
@@
There exists a randomized distributed algorithm that for any graph $G$ solves the multi-message broadcast problem in $O(\log^{3}n\cdot\tau_{mix}\cdot\mathrm{OPT})$ rounds with high probability, where $\tau_{mix}$ is the mixing time of $G$ and $\mathrm{OPT}$ is the optimal round complexity for the given problem instance.
In the analysis, we use $k/\delta(G)$ as a natural lower bound for $\mathrm{OPT}$. Indeed, if a node has degree $\delta(G)$, it requires at least $k/\delta(G)$ rounds to receive all $k$ messages. While this bound is meaningful on expander graphs, we show in Section 5.2 that it can be far from tight even on graphs with constant diameter, which highlights the necessity of exploiting expansion properties in our approach.
We design the algorithm for an arbitrary expander by building on the algorithm for random graphs, which achieves near-optimal performance in an important special case when the network is modeled as an Erdős–Rényi graph $G(n,p)$.
@@
For an Erdős–Rényi graph $G(n,p)$ with $p\geq\frac{C_{p}\log n}{n}$, there exists a distributed randomized algorithm that produces a tree packing of size $\delta(G)$, diameter $O(\log n)$, and weight $O(\log n)$ w.h.p.
@@
Finally, to map the terrain of the problem, we prove the hardness result in the centralized setting. Namely, we show that computing the exact number of rounds required for multi-message broadcast is NP-hard on general graphs.
@@ 2 Related Work
Previous Work.
The first work to address universal optimality for the multi-message broadcast problem in the CONGEST model was by Ghaffari [21]. In that paper, the algorithm consists of two phases: (1) constructing a tree packing, and (2) performing the broadcast using the constructed tree packing. With a tree packing of diameter $H$, size $S$, and weight $W$, one can complete a multi-message broadcast in $O((H+\frac{k}{S})\cdot W)$ rounds by splitting messages uniformly across the trees and propagating them sequentially within each tree. However, the limitation of [21] is that constructing the tree packing requires $\tilde{\Omega}(D+k)$ rounds, preventing the approach from achieving universal optimality. Observe that the problem can be solved in $O(D+k)$ rounds by a naive strategy: first construct a BFS tree rooted at $s$, and then downcast messages in it one by one. Thus, the result of [21] yields an improvement only when multiple consecutive instances of the problem are solved, allowing the precomputed tree packing to be reused.
@@
Routing.
One fundamental information-dissemination problem in distributed computing is routing, where the goal is to deliver a set of messages from source nodes to their respective destination nodes. Unlike broadcast, which involves sending the same message(s) to all nodes, routing requires sending individual message(s) for each source-destination pair.
@@
Unlike the above approaches, our paper provides an algorithm with only a polylogarithmic overhead.
Network Information Flow.
The network information flow problem [1] is defined as follows. The network is a directed graph $G(V,E)$ with edge capacities, a source node $s\in V$, and sink nodes $T\subseteq V$. The question is: at what maximal rate can the source send information so that all of the sinks receive that information at the same rate? In the case of a single sink $t$, the answer is given by the $\text{max-flow}(s,t)$. However, when there are multiple sinks $T$, the value $\min\limits_{t\in T}\text{max-flow}(s,t)$ may not be achievable if nodes are only allowed to relay information. In fact, the gap can be as large as a factor of $\Omega(\log n)$ [33]. Nevertheless, if intermediate nodes are allowed to send (linear [37]) codes of the information they receive, then $\min\limits_{t\in T}\text{max-flow}(s,t)$ becomes achievable [1]. Notably, in the specific case where $T=V\setminus\{s\}$ (the setting considered in the present paper), the rate of $\min\limits_{t\in T}\text{max-flow}(s,t)$ becomes achievable without coding [49]. The decentralized version of network information flow was studied in [31, 18, 30]. The most relevant work in this direction is [46] by Swamy et al., where the authors establish an optimal algorithm for the case of random graphs whose radius is almost surely bounded by $3$. Our approach works for general expanders, and in the case of random graphs, allows an expected radius to grow infinitely with $n$ (see [10] for analysis of the diameter of a random graph).
@@
Tree Packing.
The problem of tree packing has been extensively studied, as summarized in the survey by Palmer [41]. Foundational results in this area include those by Tutte [47] and Nash-Williams [40], who demonstrated that an undirected graph with edge connectivity $\lambda$ contains a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$. Edmond [14] extended this result to directed graphs, showing that such graphs always contain $\lambda$ pairwise edge-disjoint spanning trees rooted at a sender $s\in V$, where $\lambda$ is the minimum number of edges that must be removed to make some node unreachable from $s$. However, these results do not address the diameter of the tree packing.
Chuzhoy et al. [11] tackled the challenge of finding tree packings with small diameter. They presented a randomized algorithm that, given an undirected graph with edge connectivity $\lambda$ and diameter $D$, outputs with high probability a tree packing of size $\lfloor\frac{\lambda}{2}\rfloor$, weight $2$, and diameter $O((101k\log n)^{D})$.
Tree packing on random graphs was studied by Gao et al. [19], who showed that asymptotically almost surely, the size of a spanning tree packing of weight $1$ for a Erdős–Rényi graph $G(n,p)$ is $\min\left\{\delta(G),\frac{|E(G)|}{n-1}\right\}$, which corresponds to two straightforward upper bounds.
@@
Branching Random Walks in Networks.
The cover time of a random walk [36] on a graph is the time needed for a walk to visit each node at least once. Unfortunately, the expected value of this quantity is $\Omega(n\log n)$ even for a clique, making this primitive less useful in designing fast algorithms. Consequently, several attempts have been made to accelerate the cover time. Alon et al. [2] proposed initiating multiple random walks from a single source. Subsequent work by Elsässer and Sauerwald refined their bounds, demonstrating that $r$ random walks can yield a speed-up of $r$ times for many graph classes. Variations of multiple random walks have been applied in the CONGEST model to approximate the mixing time [39], perform leader election [35, 27], and evaluate network conductance [17, 4].
@@ 3 Algorithm Overview
In this section, we provide a high-level overview of our algorithm, which consists of two major parts. First, we embed a virtual Erdős–Rényi graph $G(n,p)$ atop the physical network $H$, and then we solve the problem on $G$. The reason we do this embedding is to transform an arbitrary expander into an almost-regular one. We explain the embedding procedure in Section 3.1, and from that point on, we focus solely on solving multi-message broadcast on an Erdős–Rényi graph. In Section 3.2, we overview COBRA - the main building block of our algorithm for random graphs, and in Section 3.3 we describe the algorithm itself.
@@ 3.1 Embedding a Random Graph
Our construction will guarantee that each round of communication in $G$ can be simulated in $O(\tau_{mix}(H)\cdot\log n)$ rounds in $H$ and that $\delta(G)$ will be close to $\delta(H)$. This gives us the following “lifting”: if there is an algorithm that solves a problem in $f(k,\log(|V(G)|),\delta(G))$ rounds on $G$ for some function $f$, then there is an algorithm that solves a problem in essentially $f(k,\log n,\delta(H))\cdot\tau_{mix}(H)\cdot\log n$ rounds on $H$. Formally,
@@
Assume there exists an algorithm that solves multi-message broadcast on an Erdős–Rényi graph $G$ in $O(\frac{k}{\delta(G)}\cdot\log(|V(G)|)+\log^{2}(|V(G)|))$ rounds w.h.p. Then there exists an algorithm that solves multi-message broadcast on any graph $H$ on $n$ vertices in
$O(\frac{k}{\delta(H)}\cdot\tau_{mix}(H)\cdot\log^{2}n+\tau_{mix}(H)\cdot\log^{3}n)$
rounds w.h.p.
This round complexity consists of three terms: $O((\frac{k}{\delta(H)}\cdot\log n+\log^{2}n)\cdot\tau_{mix}\cdot\log n)$ for simulating an algorithm, $O(\tau_{mix}(H)\cdot\log^{2}n)$ rounds for constructing an embedding of an Erdős–Rényi graph knowing $\tau_{mix}(H)$, and $O(\tau_{mix}(H)\cdot\log^{2}n)$ for estimating the $\tau_{mix}(H)$. Below, we describe the embedding and estimation parts.
Embedding Construction.
We now outline how to construct the embedding of an Erdős–Rényi graph atop the expander $H$ preserving the minimal degree, assuming nodes have an estimate of $\tau_{mix}$.
@@
Once paths are established, communication along $(u,v)\in E(G)$ is simulated by routing a message along the corresponding walk path in $H$, which, by Lemma 3.2, can be done in $O(\tau_{mix}(H)\cdot\log n)$ rounds.
We remark that the process described above is equivalent to sampling with replacement $\delta(H)$ neighbors for each node $v\in V(G)$ independently (almost) uniformly at random. While this is not a canonical definition of an Erdős–Rényi graph, the resulting distribution is equivalent, modulo a negligible inversely polynomial small probability.
On Distributed Estimation of Mixing Time.
Note that in order to implement this simulation, nodes do not need prior knowledge of $\tau_{mix}$. Instead, they can estimate the mixing time of the network using the decentralized algorithm of Kempe and McSherry [34], which runs in $O(\tau_{mix}\cdot\log^{2}n)$ rounds (without prior knowledge of $\tau_{mix}$). Their algorithm can be used to estimate the second principal eigenvalue $\lambda$ of the transition matrix, which relates to the mixing time via the following inequality:
@@
Given a graph on $n$ nodes with mixing time $\tau_{mix}$ and second principal eigenvalue $\lambda$, it holds that
@@
Knowing $\tau_{mix}(H)$ and $\delta(H)$, compute an embedding of an Erdős–Rényi graph $G(n^{\prime},\frac{\delta(H)}{n^{\prime}})$ as described above, where $n^{\prime}$ is the number of sub-node groups in $H$. Make the source node for a multi-message broadcast in $G$ any node in $G$ that is simulated by the source of an original problem instance in $H$. Solve the problem in $G$ in $O(\frac{k}{\delta(G)}\cdot\log n+\log^{2}n)=O(\frac{k}{\delta(H)}\cdot\log n+\log^{2}n)$ rounds, simulating each round of $G$ in $O(\tau_{mix}(H)\cdot\log n)$ rounds in $H$.